OpenStudy (anonymous):

Paul's dog was asleep 2/3 of the day. How many hours was it awake?

OpenStudy (anonymous):

1 day = 24h 24 * (2/3) = 16 --> Daul's dog slept for 16 hours. That means that the dog must've been awake for: 24 - 16 = 8 hours. @Alexis0127

OpenStudy (anonymous):

2/3 = 24/1 cross multiply and get 16, so he was asleep for 16 hours, subtract 24 from 16 and get 8 hours.

OpenStudy (anonymous):

Thank you